MIT Services Status

Core service interruptions
Incident Report for MIT
Resolved
This incident has been resolved.
Posted 27 days ago. May 22, 2018 - 12:04 EDT
Update
While the OC11 firewalls have recovered we are aware of several services that still have not recovered. We are working on stablizing those services as we learn about them.
Posted about 1 month ago. May 18, 2018 - 14:36 EDT
Monitoring
IS&T performed an emergency software upgrade to the OC11 data center firewalls. All services were restored as of 11:30am. IS&T will continue to monitor the situation.
Posted about 1 month ago. May 18, 2018 - 11:51 EDT
Investigating
IS&T is aware of another incident affecting core services and is investigating the issue.
Posted about 1 month ago. May 18, 2018 - 09:56 EDT
This incident affected: General.